OpenStudy (laddiusmaximus):

calculate y=(e^(4x))/x for y" and y"'

OpenStudy (jamesj):

Ok, so what's the first derivative, y' ?

OpenStudy (laddiusmaximus):

4e^x/1

OpenStudy (jamesj):

Use the quotient rule, write y = u/v where u = e^4x, v = x Then \[ y' = \frac{vu' - uv'}{v^2} \]

OpenStudy (laddiusmaximus):

ok yeah I knew that. crap. I did that the first time, but it didnt come out right. I had 4e^x *x - e^4x *1 / x^2

OpenStudy (laddiusmaximus):

or 4ex-4ex/2x

OpenStudy (jamesj):

Careful. It is \[ y' = \frac{x.4e^{4x} - e^{4x}}{x^2} \] \[ = \frac{4e^{4x}}{x} - \frac{e^{4x}}{x^2} \] Now differentiate that again to find y''

OpenStudy (laddiusmaximus):

so I differentiate before I use the quotient rule?

OpenStudy (jamesj):

The quotient rule is a means of differentiating the quotient of two functions. Your question doesn't make sense.

OpenStudy (laddiusmaximus):

well a lot about calculus doesnt make sense to me and my teacher stinks

OpenStudy (jamesj):

I recommend you go back in your class notes or text book and review it. You're going to need to use it quite a lot with this question.

OpenStudy (laddiusmaximus):

e^4x is just e^4x right?

OpenStudy (jamesj):

The derivative of \( f(x) = e^{ax} \) for any constant \( a \) is \[ \frac{df}{dx} = a e^{ax} \]

OpenStudy (laddiusmaximus):

because the derivative of e^x is e^x

OpenStudy (jamesj):

Yes.

OpenStudy (laddiusmaximus):

so e^4x is just e^4x, right?